Erratum: I have the same formula wrong in two different places. The formula a third of the way down on page 431 in [1] and also the formula in Proposition 6.5.6, p. 137, in [2], should be:

h_{ts} ⟼ ∑_{j=0}^{⌊ \frac{−s+t−1}{2} ⌋} ∑_{i=j+s+1}^{t−j} ζ_{j}^{2^{s}}ξ_{t−i−j}^{2^{i+j+s}} ⊗ h_{i,j+s}.

(The mistake was in the upper limit on the first sum.)

References

[1] J. H. Palmieri, Quillen stratification for the Steenrod algebra, Ann. of Math. (2) 149 (1999), 421–449.

[2] , Stable homotopy over the Steenrod algebra, Mem. Amer. Math. Soc. 151 (2001), no. 716, xiv+172.